OpenStudy (anonymous):

Find the missing length for a right triangle with hypotenuse of 18 and a leg length of 6. Round to the nearest whole number if necessary. (1 point)

OpenStudy (goformit100):



OpenStudy (goformit100):

Use Pythagoras theorem

OpenStudy (johnweldon1993):

 the pythagorean theorem states A² + B² = C² where C = hypotenuse well...here...we know the hypotenuse....so rearrange this to solve for another side A² = C² - B² plug in your numbers A² = 18² - 6² A = √(18² - 6²) what does this equal?